OpenStudy (anonymous):

find the rate if I=$900, P=$7500, and T=2 years

OpenStudy (campbell_st):

Are you expected to use I = PRT.....?

OpenStudy (anonymous):

Yes

OpenStudy (jdoe0001):

so solve for "r" then, and plug in the given values for the other elements

OpenStudy (campbell_st):

ok... so substituting you get \[900 = 7500\times R \times 2\] or \[900 = 15000 \times R\] you need to now solve for R... which will be a decimal and then change to a percentage